# Geometrically Thick Disks around Kerr Black Holes in a Swirling Universe

### v1

@@ Abstract
We investigate geometrically thick disks around Kerr black holes in a swirling universe. This stationary and axisymmetric spacetime is composed of a Kerr black hole, which is immersed in a swirling background. The swirling background is characterized by an odd $\mathcal{Z}_{2}$ symmetry, where the northern and southern hemispheres are rotating in opposite directions. The Kerr and swirling rotations can lead to the emergence of complex spin-spin interactions, which heavily influence the spacetime properties. In order to study this influence, we analyze prograde as well as retrograde circular orbits and geometrically thick disks for different spacetime solutions, which are classified by their Kerr parameter $a$ and the swirling parameter $j$. We find stabilizing effects on prograde circular orbits and destabilizing effects on retrograde circular orbits, which originate from the spin-spin interaction and depend mainly on the Kerr rotation. Furthermore, the background rotation leads to an emergence of static orbits and the symmetry breaking regarding the equatorial plane causes a vertical distribution of the circular orbits and thick torus solutions. Due to the destabilizing effect of the swirling rotation, an outer marginally stable orbit appears, which heavily downsizes the parameter space for disk solutions. Dependent on the Keplerian specific angular momentum distribution of the spacetime and on the disk momentum $\ell_{0}$, different types of disk are possible which we classify and which differ in terms of the complex disk dynamics that could arise due to the presence of an inner and outer cusp or just an outer cusp.
@@ 1 Introduction
Accretion disks are structured accumulations of matter around astrophysical compact objects. Diffuse matter gets attracted by the central object and settles in circular orbits. The formed large-scale structure resembles a disk or a torus. In complex disk interactions, angular momentum gets exchanged from the inner to the outer regions of the disk, which causes matter to migrate from the outside inwards through the disk to smaller and smaller orbits towards the central object. When the matter reaches the marginally stable orbit, small perturbations could lead to an inspiral of the matter and it ultimately gets accreted by the central object. The energy converted in this process scales with the mass of the central object and in the example of supermassive black holes, it corresponds to magnitudes that can be found at the upper end of our current energy scales for observed and theorized phenomena in the universe. Accretion disks are therefore particularly important in the study of high-energy physics and in the observation of astrophysical objects, as a large part of the energy is converted into radiation energy, with such high intensities, that it can be observed from billions of light years away. This radiation can have a broad frequency spectrum, which mainly depends on the disk properties and the properties of the central object. Disregarding various subclasses, accretion disks can generally be divided into slim and thick disks. In this paper we will focus on the class of thick accretion disks, which was first introduced in the 70s [1, 2, 3, 4, 5] and profoundly analyzed in the 80s [6, 7, 8, 9, 10].
The simplest case of a thick disk imposes the Polish Doughnut model, where the disk is modeled as a non-selfgravitating perfect fluid and where its properties are solely determined by the spacetime geometry. Due to its simplicity, this approach offers analytical solutions that are directly computed from the metric components. Nevertheless, these solutions capture qualitatively the main properties of possible disks and are therefore often used as initial conditions in more complex GRMHD simulations [11, 12, 13]. Polish Doughnuts have been studied for a variety of alternative gravity and compact object solutions [14, 15, 16, 17, 18, 19, 20, 21]. Recently Polish Doughnuts around a Schwarzschild black hole in a swirling universe were analyzed [22], where the Swirling background gives rise to unique features like an outer disk cusp. In this work, we will focus on geometrically thick tori around Kerr black holes in a swirling universe and study their properties.
In general, the swirling universe describes a rotating spacetime, which is not asymptotically flat. The formalism to obtain a swirling universe solution originates from the works of Frederick J. Ernst, who proposed in 1968 a method to obtain new solutions from the Einstein field equations for stationary and axisymmetric (electro-)vacuum spacetimes [23, 24]. In this formulation, the field equations, which are in general a system of coupled non-linear differential equations, are replaced by the Ernst equations. The Ernst equations are composed of two coupled partial differential equations for a complex gravitational and electromagnetic potential. Since these equations are integrable, they can be solved by a set of solution generating techniques, which can furthermore be applied to build new solutions from Einstein’s field equations by using the invariance of the Ernst potentials under some set of transformations. Recently, compositions of these transformations have been used to generate new solutions [25]. In particular, the swirling universe is obtained by the application of Ehlers transformation to a Minkowski seed solution and was first presented in [26]. It was rediscovered in [27], which in addition to the reintroduction of the background solution, also immersed Schwarzschild and Kerr black holes in the swirling universe, in a similar fashion like Ernst immersed black holes into magnetic universes [28]. The swirling spacetime is characterized by the swirling parameter $j$, which is responsible for the addition of exotic features to the seed solution, such as the appearance of an ergoregion that extends to infinity, in the presence or absence of compact objects. A detailed description of that ergoregion and the geodesic motion in the swirling universe can be found in [29]. The swirling spacetime is stationary, axisymmetric, and harbours an odd $\mathcal{Z}_{2}$ symmetry, where the background frame dragging changes its sign in the lower hemisphere (with respect to the equatorial plane). As a consequence, the spacetime in the lower hemisphere rotates in the opposite direction as in the upper hemisphere. The spacetime of a Schwarzschild black hole immersed in this swirling universe (SBHSU) is fully characterized by the mass of the black hole and by the swirling parameter of the spacetime. The spacetime of a Kerr black hole immersed in a swirling universe (KBHSU), is fully characterized by the black hole mass, the Kerr parameter, and the swirling parameter. The KBHSU is co-rotating with the swirling background in one hemisphere and counter-rotating in the other since the Kerr parameter does not change its sign with respect to the equatorial plane. This breaking of symmetry and the spin-spin interaction of the Kerr rotation with the swirling rotation has an influence on the main properties of the spacetime, like the ergoregions, as well as the circular orbits and accretion tori solutions, which are usually symmetrical with respect to the equatorial plane. The main goal of this work is to analyze these properties, how they are affected by the swirling parameter for different Kerr parameters, and how they change with variation of the swirling and Kerr parameters.
@@
The structure of this work is as follows: In section 2 we introduce the KBHSU solution and briefly discuss its main properties. In section 3 a introduction to the theory of geometrically thick accretion disks is presented, and in section 4 we analyze qualitatively the circular orbits and thick disk solutions of the KBHSU spacetime. Throughout this work we use the Einstein convention and geometrized units, where $G=c=1$, with $G$ as the gravitational constant and $c$ as the speed of light. The metric signature used is (,+,+,+)(-,+,+,+)( - , + , + , + ) and in all KBHSU solutions throughout this work, the mass of the KBHSU is normalized to $M=1$.
2 Kerr black hole in a swirling universe
@@
The line element describing a Kerr black hole immersed in a swirling universe can be written in Boyer-Lindquist coordinates as [27]
@@
${\rm d}s^{2}=\frac{1}{\mathcal{F}(r,\theta)}\left({\rm d}\varphi+\omega{\rm d}t\right)^{2}+\mathcal{F}(r,\theta)\left[-\rho^{2}{\rm d}t^{2}+\Sigma\sin^{2}\theta\left(\frac{{\rm d}r^{2}}{\Delta}+{\rm d}\theta^{2}\right)\right],$ (1)
@@
$\Delta$ $=$ $r^{2}-2mr+a^{2},\ \ \ \qquad\ \ \ \rho^{2}=\Delta\sin^{2}\theta,\ \ \ \qquad\ \ \ \Sigma=\left(r^{2}+a^{2}\right)^{2}-\Delta a^{2}\sin^{2}\theta,$
@@
The pure Kerr solution is recovered for vanishing $j$, the immersion of Schwarzschild in a swirling universe is obtained by $a=0$, and the pure swirling universe is found for $m=0$ and $a=0$. As noted in [27] the spin-spin interaction between the Kerr black hole and the background dragging gives rise to a conical singularity on the symmetry axes. Furthermore, the KBHSU solution possesses two horizon surfaces, an inner and an outer one, which in these coordinates are found by imposing $g^{rr}=0$, and thus are defined by
@@
which is independent of $j$ and equivalent to the Kerr solution. However, it should be noted, that the presence of the background dragging will deform the horizon geometry. Considering the main interest of this paper, we will focus only on the outer horizon here. Taking a slice of constant $t$ leads to the induced metric
@@
Insights about the shape of the horizon can be taken by considering an isometric embedding of the two-dimensional surface (4) into a three-dimensional Euclidean space. However, a full embedding is typically not always possible. For the pure Kerr solution, i.e. $j=0$, the event horizon geometry around the poles becomes hyperbolic as the Kerr parameter increases ($a\in(\sqrt{3}/2,1)$) and thus an embedding of those regions into a Euclidean space is not feasible. For that case, insights about the geometry can be shown by considering the composition of a Euclidean embedding with the metric $ds^{2}_{E}=dx^{2}+dy^{2}+dz^{2}$ and a pseudo-Euclidean embedding with metric $ds^{2}_{pE}=dx^{2}+dy^{2}-dz^{2}$ [30, 31].
@@
The presence of the swirling background affects the north and south hemispheres differently, which is also directly reflected in the horizon geometry. We note that even small variations of $j$ deform the usual oblate structure of the Kerr horizon. For a positive spin-spin relation ($aj>0$) the neighborhood of $(r=r_{+},\theta=\pi)$ becomes hyperbolic. For a rapidly rotating Kerr black hole, where the neighborhood of the poles is already hyperbolic, the vicinity of $(r=r_{+},\theta=0)$ is dragged into an elliptic geometry. In Fig. 1 we show a representation of the outer horizon for a variation of the swirling parameter, $j\in\{0,10^{-4},10^{-3}\}$, and two values of the Kerr parameter, $a\in\{0.5,0.9\}$. For a negative spin-spin relation ($aj<0$) an analogous behaviour would be observed, where just the northern and southern hemispheres are interchanged.
@@
Figure 1: Embeddings of the outer horizon geometry for different KBHSU solutions and small variations of the swirling parameter $j$. The orange curves and surfaces represent an Euclidean embedding, the cyan curves and surfaces represent a pseudo-Euclidean embedding. The top row shows the horizon geometry for a KBHSU with Kerr parameter $a=0.5$, where the vicinity of the southern pole becomes hyperbolic due to the swirling dragging. This effect increases with $j$. The middle row shows the geometry for a rapidly rotating KBHSU with $a=0.9$. Here, the regions close to the northern and southern poles are already hyperbolic and with increasing $j$ the horizon region close to the northern pole becomes elliptic. The bottom row presents a 3D representation of the embedding for the rapidly rotating case with $a=0.9$. The equatorial plane has been set to be $Z(\pi/2)=0$ in all plots.
When $j$ increases further, the geometry of the event horizon is significantly modified. Some examples are shown in Fig. 2, where we have chosen the limiting Kerr parameter $a=\sqrt{3}/2$ to illustrate the hyperbolic behaviour appearing due to the swirling dragging.
@@
Despite the significant changes in geometry, the horizon area is not affected by the presence of the swirling parameter $j$,
which has already been seen for the Schwarzschild black hole in a swirling universe [27, Moreira:2024sjq].
The ergosurfaces for the KBHSU solution are rather complicated and change significantly depending on $j$. Since it appears not to be possible to find simple analytical relations, as it is possible for pure Kerr and the swirling universe background, we perform a qualitative study on the boundaries of the ergoregions, which are defined by the condition $g_{tt}=0$. The pure Kerr solution has two of these ergosurfaces, an inner and an outer closed surface that touches the respective horizons on the symmetry axis, thus
$r_{E_{-}}(m,a,\theta)\leq r_{-}<r_{+}\leq r_{E_{+}}(m,a,\theta),$ (6)
where $r_{E_{\pm}}$ is the radial coordinate of the inner and outer ergosurface, respectively. These surfaces are also present in the KBHSU solution and the constraints given in Eq. 6 still hold for the KBHSU case. Furthermore, two additional ergosurfaces appear, which for small values of $j$ are disconnected patches in the southern and northern hemispheres extending to infinity. These surfaces are a property of the swirling background. They appear also for the Schwarzschild black hole immersed in a swirling universe. A detailed discussion of these ergosurfaces can be found in [27, 29, Moreira:2024sjq]. For the main interest of this paper, we will restrict the discussion of the ergoregions to the outer area, i.e. $r>r_{+}$.
The ergoregions of the KBHSU are intrinsically related to the spin-spin interaction of black hole and swirling background. For smaller variations of $j$, the ergosurfaces resemble roughly a superposition of the intrinsic ergoregions from each spacetime (Kerr and swirling universe) described above. This is illustrated in Fig. 3 below. In Fig 4 we show a representation of these surfaces in a three-dimensional space for a small value of $j=0.01$.
@@
Figure 3: Comparison of the ergosurfaces for KBSHU, Kerr, and the swirling universe (SU) in the $r-\theta$ plane for two different values of the swirling parameter. The orange curves show the outer KBHSU ergosurfaces, the blue dashed curve shows the outer Kerr ergosurface, and the red dashed curves show the swirling universe ergosurfaces. The Kerr parameter is set to a $a=0.9$ in both plots.
The KBHSU ergosurface, which corresponds to the outer Kerr surface remains qualitatively similar for small values of $j$ (Fig. 3 (a)). The ergosurfaces originating from the swirling background show a minor deviation for smaller $j$, but their general structure is similar. However, as $j$ increases, the ergosurfaces deviate more and more from the Kerr and swirling universe surfaces and are not just a combination of these any longer. The non-linear spin-spin interaction causes greater symmetry breaking for larger $j$ and the ergosurfaces become more asymmetrical with respect to the equatorial plane.
Figure 4: Representation of the outer KBHSU surfaces in a three-dimensional space with $Z(\pi/2)=0$. The Kerr parameter is set to $a=0.9$ and the swirling parameter to $j=0.01$. The inner structure showcases the ergosurface which originates from the outer Kerr surface and the structures in the northern and southern hemispheres showcase the ergosurfaces, which originate from the swirling universe ergosurfaces. This representation is obtained by considering the rotation around the symmetry axis.
Generally, we conclude, that the spin-spin interaction between the black hole and the swirling background enhances the symmetry breaking regarding the spacetime properties, as showcased for the horizons and ergosurfaces. This symmetry breaking also has an influence on the properties of circular orbits and thick disks as we will showcase in the following sections. For further reading on the general properties of the KBHSU spacetime, we can point out, that a more detailed discussion is in development.

### v2

# Accretion Structures around Kerr Black Holes in a Swirling Background

@@ Abstract
In this paper, we investigate thick accretion structures around a Kerr black hole in a swirling background, that is, a rotating black hole immersed in a rotating background. This is a novel solution characterized by the black hole mass in addition to two distinct rotational parameters, the Kerr parameter $a$, which identifies the rotation of the black hole, and the swirling parameter $j$, which describes the background rotation. The swirling background is characterized by an odd $\mathcal{Z}_{2}$ symmetry, where the northern and southern hemispheres rotate in opposite directions. The rotation of the black hole embedded into this swirling background leads to non-trivial spin-spin interactions with the background rotation. The spacetime properties in the vicinity of the black hole are significantly influenced by this spin-spin interaction. In order to study the influence on the basic properties of this spacetime, we analyze circular orbits and geometrically thick disks for different spacetime solutions, which are classified by the black hole and swirling spins. We identify stabilizing effects on prograde circular orbits and destabilizing effects on retrograde circular orbits, which originate from the spin-spin interaction and depend mainly on the Kerr rotation. Furthermore, we discovered the emergence of static orbits, which appear due to the background rotation. The symmetry breaking of the spacetime rotation with regard to the equatorial plane highly influences the spatial distribution of circular orbits. This asymmetry causes a concave (convex) distribution of the prograde (retrograde) circular orbits and accordingly, bowl-like deformations of the accretion disk solutions. Moreover, due to the destabilizing effect of the swirling rotation, an outer marginally stable orbit appears, which heavily limits the range of the parameter space in which disk solutions can exist. Due to the possibility of an outer and inner disk cusp, different types of disk solutions are possible. We classify the different types of disk solutions, which differ from each other by the properties of their cusps. Four different scenarios can be identified in which different accretion dynamics could arise.
@@ 1 Introduction
In the last century, the paradigm of gravitational physics has been significantly modified by the publication and development of general relativity. Among its most extraordinary predictions is the existence of black holes, which are believed to be the final stage of a stellar collapse. Obtaining solutions from the field equations of general relativity, i.e., Einstein’s field equations, has shown to be a herculean task. Of particular interest is the existence of exact solutions, since those classes of solutions allow to shed light on the main properties of gravity using analytical techniques. Most prominent are the black hole solutions of the coupled (electro-)vacuum Einstein(-Maxwell) field equations. In general, physically meaningful solutions are found by imposing symmetries on the field equations. Among the most well-known solutions is the Schwarzschild black hole, which describes the exterior gravitational field of a spherical compact object; the rotating counterpart of such solution is the Kerr black hole, uniquely described by the black hole mass and angular momentum. However, the simplest solution possessing rotation must be stationary and axially symmetric. The general formulation to classify solutions with such symmetries was developed by Ernst [1, 2], in which the coupled Einstein-Maxwell field equations are replaced by the Ernst potentials. Besides simplifying the study of the field equations under the assumed symmetries, this formalism allows to derive new solutions from old known (“seed”) solutions by exploring hidden symmetries within the Ernst potentials, which is often referred to as Ernst generating techniques [3, 4]. In particular, two transformations are often used to generate new solutions: a) the Harrison transformation, which allows the embedding of black holes into a magnetic universe [3, 5], and b) the Ehlers transformations, which generates the swirling background and the immersion of compact objects into it [6]. The latter is a solution that, although it has appeared before [7], did not receive much attention until it was rediscovered more recently in [6], where the physical properties of this novel family of solutions have been extensively studied. Overall, the swirling background is a non-asymptotically flat solution that describes a rotating background equipped with an odd $\mathcal{Z}_{2}$ symmetry. The spacetime is characterized by the solution’s frame-dragging, which has an opposite sign in the different hemispheres; thus, the spacetime rotates in opposite directions below and above the equatorial plane. This solution is uniquely characterized by just one parameter, the swirling parameter $j$, which gives rise to the presence of exotic features, such as the appearance of an ergoregion that extends to infinity. A detailed description of the ergoregion and the geodesic motion in the swirling background can be found in [8]. Despite these odd characteristics, this novel solution does not possess other exotic features that often appear in rotating background spacetimes, like for the Gödel universe or for the Taub-NUT spacetime, which contain closed timelike curves [9, 10]. Due to the similarity in the derivation of the swirling family of solutions, it has been taken to as a cousin solution to the magnetic universe. Moreover, it has been pointed out that such rotating spacetime could be considered locally to model cosmic filaments [11], which are believed to be the largest structures in our universe. They are made of galaxy clusters arranged in thread-like structures [12]. Interestingly, observations reveal that such structures do rotate [13, 14] and can even possess electromagnetic fields [15]. Thus, the swirling background could potentially be used to model such cosmic filaments. Therefore, the effects of such background on dispersed matter, especially around compact objects immersed in such non-trivial geometry, posts itself as a relevant topic.
@@
In the vicinity of compact objects, diffuse matter gets attracted by the gravitational pull of the compact object. Over longer time periods, accumulations of matter orbiting the compact object gradually form accretion structures. Since these structures usually have the geometry of a disk or a torus, they are called accretion disks or tori and are characterized by the circular flow of their disk particles. In complex disk interactions, angular momentum gets exchanged from the inner to the outer regions of the disk, which causes matter to migrate from the outside inwards through the disk to smaller and smaller orbits towards the central object. When matter reaches the marginally stable orbit, small perturbations could lead to an inspiral of the matter onto the central object, where it ultimately gets accreted. The energy converted in this process scales with the mass of the central object, and in the case of supermassive black holes, it corresponds to magnitudes that can be found at the upper end of our current energy scales for observed and theorized phenomena in the universe. Accretion disks are therefore particularly important in the study of high-energy physics and in the observation of astrophysical objects, as a large part of the energy is converted into radiation energy, with such high intensities that it can be observed from billions of light years away. This radiation can have a broad frequency spectrum, which mainly depends on the properties of the disk as well as on the properties of the central object. Overall, multiple approaches exist in the modeling of accretion disks, which differ by the dynamical situations they are intended to describe. When it comes to their morphology, they can generally be divided into two classes with regard to their vertical height, namely, thin and thick disks. In this paper, we focus on the latter, whose theoretical foundations have been vastly established in [16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
The simplest case of a thick disk imposes the Polish Doughnut model, in which the disk is modeled as a body composed of a non-selfgravitating perfect fluid, whose properties are solely determined by the spacetime geometry. Due to its simplicity, this approach often results in analytical solutions that can be directly computed from the metric tensor. Nevertheless, despite its simplicity, such modeling captures the main qualitative properties of possible accretion structures and is often used as an initial condition in more complex general relativistic magnetohydrodynamics simulations [26, 27, 28]. The Polish Doughnut model has been applied to a variety of alternative gravity theories and compact objects, such as de Sitter black holes, scalarized black holes, NUT spacetimes, and Born-Infeld teleparallel gravity [29, 30, 31, 32, 33, 34, 35, 36]. Furthermore, the thick disk model has recently been generalized to the swirling family of solutions, more precisely to the Schwarzschild black hole immersed in a swirling background [37]. The spacetime of a Schwarzschild black hole immersed in a swirling background is fully characterized by the mass of the black hole and the swirling parameter $j$. It was shown that the presence of the rotating background modifies the typical disk structures around the Schwarzschild black hole by deforming the thick disk into a bowl-shaped matter distribution as well as by the appearance of an outer disk cusp, which originates from instabilities of the circular geodesics within the spacetime [37].
Here, we will focus on generalizing this work to the Kerr black hole immersed in a swirling background (KBHSB). This spacetime is characterized by three parameters, namely the black hole mass $M$, the Kerr spin parameter $a=J/M$, where $J$ is the angular momentum of the black hole, and the swirling parameter $j$. In contrast to the swirling motion, the Kerr rotation does not change its sign with respect to the equatorial plane; therefore the black hole is co-rotating with the swirling background in one hemisphere and counter-rotating in the other. This breaking of symmetry and the resulting spin-spin interaction of the Kerr rotation with the swirling rotation has an influence on the main properties of the local spacetime environment. The motion of particles in the local vicinity of the black hole is especially affected by the spin-spin interaction and the symmetry breaking. As such, unique effects arise on the properties of basic geodesic motion around the black hole. In this work, we aim to shed light on these effects by analyzing the influence of the Kerr and swirling rotations on the properties of circular orbits and geometrically thick accretion structures.
@@
This work is organized as follows. In section 2, we introduce the KBHSB solution and briefly discuss its main characteristics. In section 3, an introduction to the theory of geometrically thick accretion disks is presented, and in section 4 we analyze qualitatively the circular orbits and thick disk solutions of the KBHSB spacetime. Throughout this work, we use the Einstein summation convention and geometrized units, where $G=c=1$, with $G$ as the gravitational constant and $c$ as the speed of light. Note that, in this units, the swirling parameter $j$ has the dimension $[j]=1/M^{2}$. The metric signature used is (,+,+,+)(-,+,+,+)( - , + , + , + ) and for all KBHSB solutions considered in this work, the black hole mass is normalized to $M=1$.
2 Kerr black hole in a swirling background
@@
The line element describing a Kerr black hole immersed in a swirling background can be written in Boyer-Lindquist coordinates as [6]
@@
${\rm d}s^{2}=\frac{1}{\mathcal{F}(r,\theta)}\left({\rm d}\varphi+\omega{\rm d}t\right)^{2}+\mathcal{F}(r,\theta)\left[-\varrho^{2}{\rm d}t^{2}+\Sigma\sin^{2}\theta\left(\frac{{\rm d}r^{2}}{\Delta}+{\rm d}\theta^{2}\right)\right],$ (1)
@@
$\Delta$ $=$ $r^{2}-2mr+a^{2},\ \ \ \qquad\ \ \ \varrho^{2}=\Delta\sin^{2}\theta,\ \ \ \qquad\ \ \ \Sigma=\left(r^{2}+a^{2}\right)^{2}-\Delta a^{2}\sin^{2}\theta,$
@@
The pure Kerr solution is recovered for vanishing $j$, the immersion of Schwarzschild in a swirling background is obtained by $a=0$, and the pure swirling background is found for $m=0$ and $a=0$. As noted in [6] the spin-spin interaction between the Kerr black hole and the background frame dragging gives rise to a conical singularity on the symmetry axis. The presence of conical singularities in the immersion of different spacetimes into the swirling background was already pointed out in [38], where the accelerated Kerr solution was considered and the metric could be regularized by fixing the swirling parameter. Nonetheless, this appears not to be possible in the KBHSB case. Furthermore, the KBHSB solution possesses two horizon surfaces, an inner and an outer one, which in these coordinates are found by imposing $g^{rr}=0$, and thus are defined by
@@
which is independent of $j$ and equivalent to the Kerr solution. However, it should be noted that the presence of the background dragging does deform the horizon geometry. Considering the main interest of this paper, we will focus only on the outer horizon here. Taking a slice of constant $t$ leads to the induced metric
@@
Insights about the shape of the horizon can be taken by considering an isometric embedding of the two-dimensional surface (4) into a three-dimensional Euclidean space. However, a full embedding is not always possible. For the pure Kerr solution, i.e. $j=0$, the event horizon geometry around the poles becomes hyperbolic as the Kerr parameter increases ($a\in(\sqrt{3}/2,1)$), and the embedding of those regions into Euclidean space is not feasible. For this purpose, insights about the geometry can be shown by considering the composition of an Euclidean embedding with the metric $ds^{2}_{E}=dx^{2}+dy^{2}+dz^{2}$ and a pseudo-Euclidean embedding described by the metric $ds^{2}_{pE}=dx^{2}+dy^{2}-dz^{2}$ [39, 40].
@@
The presence of the swirling background affects the north and south hemispheres differently, which is also directly reflected by the horizon geometry. We note that even small variations of $j$ deform the usual oblate structure of the Kerr horizon. For a positive spin-spin relation ($aj>0$) the neighborhood of $(r=r_{+},\theta=\pi)$ becomes hyperbolic. For a rapidly rotating Kerr black hole, where the neighborhood of the poles is already hyperbolic, the vicinity of $(r=r_{+},\theta=0)$ is dragged into an elliptic geometry. In Fig. 1 we show a representation of the outer horizon for a variation of the swirling parameter $j\in\{0,10^{-4},10^{-3}\}$, and the Kerr parameter, $a\in\{0.5,0.9\}$. For a negative spin-spin relation ($aj<0$) an analogous behavior would be present, where just the northern and southern hemispheres are interchanged.
When $j$ increases further, the geometry of the event horizon is significantly modified. Some examples are shown in Fig. 2, where we have chosen the limiting Kerr parameter $a=\sqrt{3}/2$, for which the horizon geometry starts to become hyperbolic, to illustrate the hyperbolic behavior appearing due to the swirling dragging.
Despite the significant changes in geometry, the horizon area $\mathcal{A}$ is not affected by the presence of the swirling parameter $j$,
which has already been seen for the Schwarzschild black hole in a swirling background [6, 41]. The ergosurfaces for the KBHSB solution are rather complicated and change significantly depending on $j$. Since it appears not to be possible to find simple analytical relations, as it is possible for pure Kerr and the swirling background, we perform a qualitative study on the boundaries of the ergoregions, which are defined by the condition $g_{tt}=0$. The pure Kerr solution has two of these ergosurfaces, an inner and an outer closed surface that touches the respective horizons on the symmetry axis, thus
$r_{E_{-}}(m,a,\theta)\leq r_{-}<r_{+}\leq r_{E_{+}}(m,a,\theta),$ (6)
where $r_{E_{\pm}}$ is the radial coordinate of the inner and outer ergosurface, respectively. These surfaces are also present in the KBHSB solution and the constraints given in Eq. 6 still hold for the KBHSB case. Furthermore, two additional ergosurfaces appear, which, for small values of $j$, are disconnected patches in the southern and northern hemispheres extending to infinity. These surfaces are a property of the swirling background. They also appear for the Schwarzschild black hole immersed in a swirling background. A detailed discussion of these ergosurfaces can be found in [6, 8, 41]. For the main interest of this paper, we will restrict the discussion of the ergoregions to the outer area, i.e. $r>r_{+}$.
The ergoregions of the KBHSB are intrinsically related to the spin-spin interaction between the black hole and swirling background. For smaller variations of $j$, the ergosurfaces resemble roughly a superposition of the intrinsic ergoregions from each spacetime (Kerr and swirling background) described above. This is illustrated in Fig. 3. In Fig. 4, we show a representation of these surfaces in a three-dimensional space for a small swirling parameter value of $j=0.01$.
The KBHSB ergosurface, corresponding to the outer Kerr surface, remains qualitatively similar for small values of $j$ (Fig. 3 (a)). The ergosurfaces originating from the swirling background show a minor deviation for smaller $j$, but their general structure is similar. However, as $j$ increases the ergosurfaces deviate more and more from the Kerr and swirling background surfaces and are no longer resembling a superposition of the two limiting solutions. The non-linear spin-spin interaction causes greater symmetry breaking for larger $j$, and the ergosurfaces become more asymmetrical with respect to the equatorial plane.
Generally, we conclude that the spin-spin interaction between the black hole and the swirling background enhances the symmetry breaking with regard to the major spacetime properties, as showcased for the horizons and ergosurfaces. This symmetry breaking also has an influence on the properties of circular orbits and thick disks, as it will be demonstrated in the following sections.
@@
Figure 1: Embeddings of the outer horizon geometry for different KBHSB solutions and small variations of the swirling parameter $j$. The orange curves and surfaces represent an Euclidean embedding, the cyan curves and surfaces represent a pseudo-Euclidean embedding. The top row shows the horizon geometry for a KBHSB with Kerr parameter $a=0.5$, where the vicinity of the southern pole becomes hyperbolic due to the swirling dragging. This effect increases with $j$. The middle row shows the geometry for a rapidly rotating KBHSB with $a=0.9$. Here, the regions close to the northern and southern poles are already hyperbolic and with increasing $j$ the horizon region close to the northern pole becomes elliptic. The bottom row presents a 3D representation of the embedding for the rapidly rotating case with $a=0.9$. The equatorial plane has been set to be $Z(\pi/2)=0$ in all plots.
@@
Figure 3: Comparison of the ergosurfaces for KBHSB, Kerr, and the swirling background in the $r-\theta$ plane for two different values of the swirling parameter. The orange curves show the outer KBHSB ergosurfaces, the blue-dashed curve shows the outer Kerr ergosurface and the red-dashed curves show the swirling background ergosurfaces. The Kerr parameter is set to $a=0.9$ in both plots.
Figure 4: Representation of the outer KBHSB surfaces in a three-dimensional space with $Z(\pi/2)=0$. The Kerr parameter is set to $a=0.9$ and the swirling parameter to $j=0.01$. The inner structure showcases the ergosurface originating from the outer Kerr surface, and the structures in the northern and southern hemispheres showcase the ergosurfaces from the swirling background. This representation is obtained by conducting a rotation around the symmetry axis.